OpenStudy (anonymous):

COMPLEX NUMBERS? - If (a+ib)^3 =8 , prove that a^2 + b^2 =4.

OpenStudy (primeralph):

Try expanding first.

OpenStudy (anonymous):

THE HINT SAID WE SHOULD EXPRESS b^2 in terms of a, and then solve for a....do not know what that means \[b ^{2} = 4-a ^{2}\]

OpenStudy (primeralph):

I just told you where to start.

OpenStudy (anonymous):

ok, let me write what I found after expansion

OpenStudy (anonymous):

\[a ^{3}+3a ^{2}ib-3ab ^{2}-ib ^{3}\]

OpenStudy (primeralph):

Since the result is 8, the imaginary parts have to be equal to 0 because 8 is the same as 8+0i.

OpenStudy (primeralph):

Carry on.

OpenStudy (anonymous):

so \[a ^{3}-3ab ^{2} =8\] \[3a ^{2}b-b ^{3}=0\]

OpenStudy (primeralph):

Yeah, and so on.

OpenStudy (anonymous):

i want to assume that i should then solve for a^2 and b^2 and then substitute into \[a ^{2}-b ^{2}=4\] ?

OpenStudy (primeralph):

Thought it was a +.

OpenStudy (anonymous):

oh yes +

OpenStudy (primeralph):

You can solve it however way you want.

OpenStudy (anonymous):

thank u!

OpenStudy (primeralph):

That method might be hard though. Just try manipulating the equations.

OpenStudy (anonymous):

i can see that here, i am getting nowwhere

OpenStudy (primeralph):

I'm actually getting 4/a, not just 4.

OpenStudy (primeralph):

Ignore that.

OpenStudy (anonymous):

it was an exam q , as is

OpenStudy (primeralph):

So basically factor out the part that equals 0. You get that b=0 or 3a^2-b^2 = 0. If b=0, then the other equation reduces to a^3 = 8. So a = 2 and a^2+b+2 = 4. Now for the second case where 3a^2-b^2 = 0....

OpenStudy (anonymous):

ok let me continue from there...

OpenStudy (anonymous):

\[2^{2}-0^{2}=4\]

OpenStudy (primeralph):

Eventually, you'll get another pair of solutions: a = -1 and b = sqrt3.

OpenStudy (anonymous):

where is this one coming from?

OpenStudy (primeralph):

There are always two solutions to a quad equation.

OpenStudy (primeralph):